- New and returning student employees will need refresher training about shelving and shelf reading, including the recording of statistics
- Shift at Kilmer is now complete, despite delays because of the unit being closed a week, temporary range guides are in place until permanents arrive.
- An estimated 102,000 items in Alex still need IRIS records or additions to existing records. New recon process will be to add records into IRIS without verifying and bar-coding the items on the shelf. There may be approx. 10,200 that are not on the shelf and may end up getting charged to missing. Project will commence soon and should take about two years. The group recommended an accelerated missing / withdrawal process for these items (easily recognized by the auto generated SIRSI barcode numbers
- The subcommittee met with Farideh to discuss the groups’ proposals for food and drink policies and public awareness campaign.
- Display and posters should be in a professional style, which reflects the image of a serious academic environment, and highlight the fact that collection can not be replaced or repaired so it should be protected.
- revisions are complete and final copy will be given to Roger to post online.
- so far 17 people have contacted Nick to express interest in attending a training session
- Once details have been solidified, he will send notification about the dates and times of the sessions
ANDY/MATH- continues to inventory collection which is about half done and has found a number of books that are missing, not found in catalog, or have call number errors, problems are forwarded to selectors or database management.
NICK/LSM- preparing for the Gov Docs training sessions, which may take place at two locations to for the convenience of staff on different campuses.
EDDIE/LSM- the shift is progressing and is now up to “F’s”, space for future growth of several sections was decreased to accommodate 3 years rather than 5 to make the necessary space.
EUGENE/SMLR- an archivist was hired for one year to weed and relocate items in the building’s basement, new director is getting acclimated and her presence is appreciated.
NICOLE/CHANG- busy processing fall reserve materials, continuing work on MARC holdings conversion project.
MATT/DOUGLASS- shift of “H’” will be done by the end of the month then a shift of “GV’s” will begin, familiarizing new student employees with shelving and shelf reading procedures.
ROGER/LSM- working with Marc Forester to train new student employees, helping plan a follow-up session for those involved in MARC holdings project, which will include more streamlined procedures.
ERIKA/ART- busy training new student employees, preparing for an upcoming exhibit on memorial art work and encourages all to attend.
MELODY/LSM- working with Roger on follow-up training session for those involved in the MARC holdings conversion project, and training new student employees
RENEE/LSM- busy shelving periodicals, and working on creating an area for a new book collection.
MEGAN I. /PHYSICS- training new student employees, working on MARC holdings conversion project.
MARY B./ALEX – working on a project that involves tracking of books charged to CS-ALEX and creating guidelines for use of CS-LSM/CS-ALEX.
SHIRLEY/ALEX – noticing a dramatic decrease in IN-PROCESS books, but an increase in receipt of items that are in very poor and unsanitary condition, which have possibly been donated.
ROB/ALEX- working on items for the preservation display subcommittee, shelving in REF has been braced and weeding has begun, noticing a significant ant problem on several floors.
IRINA/ALEX- continues to attend bindery repair training, processing reserve material for fall semester.
JEFF/ALEX- working on developing standardized guidelines for use of CS-LSM/CS-ALEX, also involved in the project to reduce number of circ on the fly books.
IAN/NBL- freeze on binding continues and doesn’t know when it will be lifted, continues to lead bindery repair training sessions, will start a project to insert loose pages from missing from bound periodicals at LSM.
JAMIE/LSM- working on processing items from Chemistry’s weeding project, and repairing damaged material.
Next meeting will be October 8, 2003
2:30 p.m. at SMLR, Rm 115
All future meetings will be held at 2:30 p.m.